OpenStudy (anonymous):

what's the 1st and 2nd derivative for y= (1+2x^2)/(1+x^2) what's the 1st and 2nd derivative for y= (1+2x^2)/(1+x^2) @Mathematics

OpenStudy (anonymous):

i got 2x/(1+x^2)^2 for the first one, but i know it has to be wrong bc my graph isn't turning out right

OpenStudy (anonymous):

bc?

OpenStudy (agreene):

\[\frac{dy}{dx}\frac{1+2x^2}{1+x^2}=\frac{2x}{(x^2+1)^2}\] use quotient rule for y' \[\frac{dy}{dx}\frac{2x}{(x^2+1)^2}=\frac{2-6x^2}{(x^2+1)^3}\] for y" use product rule and then chain rule what's left.

OpenStudy (anonymous):

give me a medal

OpenStudy (anonymous):

post areply and i give you

hero (hero):

Cindee, did you input it as 2x/(1+x^2)^2

hero (hero):

If you input it linearly, you have to include the parentheses

OpenStudy (anonymous):

yes i did, it had to sketch the graph by hand. I compared it with the calculator after and it didn't match up. I think i messed up on the algebra when doing to second derivative. Is there any easier or faster why i getting the second derivative without using the quotient rule?

hero (hero):

Well, if there's a sqrt in the denominator you can use conjugation to get rid of the sqrt in the denominator, but that's about it. You really can't avoid using quotient rule

OpenStudy (anonymous):

sigh* ok it just gets so messy sometimes. Thank you!!

OpenStudy (agreene):

I didn't use the quotient rule on the 2nd derivative.

hero (hero):

Do yourself a favor and get a TI-Nspire CAS

OpenStudy (anonymous):

we aren't allowed to use calculators for exams..

hero (hero):

A shame

OpenStudy (agreene):

like I said in my post: "for y" use product rule and then chain rule what's left" I prefer TI89 or TI92 over the Nspire series, for no real reason... lol

OpenStudy (anonymous):

@agreene how did you move the denominator to the numerator?

OpenStudy (agreene):

I didnt

hero (hero):

TI-Nspire is excellent if you know how to use it

hero (hero):

I know why you prefer the old OS types

hero (hero):

Quicker input

OpenStudy (agreene):

'Cause I'm poor mostly, lol

hero (hero):

Well, the 1st Generation TI-Nspires are only about $50 bucks now

OpenStudy (anonymous):

i'm confused then how did you use the product rule to get the second derivative

OpenStudy (agreene):

factor out the 2 and get: 2(d/dx [ x/((x^2+1)^2)] product rule is: d/dx (u*v) = v du/dx + u dv/dx take u = x and v= 1/((x^2+1)^2) then you are at: \[2(x(\frac{d}{dx}(\frac{1}{(x^2+1)^2}))+\frac{\frac{d}{dx}(x)}{(x^2+1)^2}\]

OpenStudy (agreene):

Then you can chain rule and start reducing.